An object with a net charge of 40 µC is placed in a uniform electric field of 629 N/C, directed vertically. What is the mass of the object if it floats in the electric field? kg
Answer #1

Since , object is floating,

So, gravitational force on object(Fg) = electric force on object(Fe)

m*g = q*E

here, m =mass of object = ??

g = 9.81 m/sec^2

q = 40 \muC

E = 629 N/C

So, m = (q*E)/g

m = (40*10^-6)*(629)/(9.81)

m = 2.56*10^-3 kg

Please upvote.